OpenStudy (anonymous):

Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b^2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value, given the following quadratic equation: 2x^2 + 7x + 3 = 0 Numerical Answers Expected!

OpenStudy (campbell_st):

b^2 - 4ac if its positive and a perfect square, then the quadratic can be factored

OpenStudy (campbell_st):

so b^2 - 4ac = 49 - 24 = 25 positive and a perfect square so the quadratic can be factored (2x + 1)(x + 3)